# Deterministic Scheduling over Wi-Fi 6 using Target Wake Time: An Experimental Approach

### v1

@@ I-B Wi-Fi 6
I-C CSMA/CA
Carrier-Sense Multiple Access with Collision Avoidance (CSMA/CA) is a channel access method traditionally used by Wi-Fi devices. A CSMA/CA client listens to check for an idle channel before transmission. If the channel is indeed busy, it waits till the end of the transmission plus the Distributed Inter Frame Space (DIFS) duration before choosing a random delay. This delay is uniformly chosen from the contention window. If any other client transmits during this period, the countdown is paused and resumed after the transmission is completed. In case, there is a collision, the client uses a binary exponential back-off mechanism before trying again, wherein the contention window is doubled and a new delay is sampled from it. The back-off, on average, doubles after every failure, thus reducing the collision probability.
However, the CSMA/CA based channel access may cause significant penalties in terms of delay and throughput, while reducing the overall efficiency of the system due to airtime wasted as a result of the back-off mechanism. The CSMA/CA also impacts the fairness of the transmission. In a scenario where the traffic is downlink biased, yet a few clients with uplink traffic may receive a higher proportion of transmit time. This is due to the fact that over a period of time the odds of every device winning the channel average out to be the same and when the Access Point (AP) wins the channel, its transmit time has to be further divided among multiple downlink clients. This is often referred to as the uplink/downlink disparity of probabilistic channel access [4].
The CSMA/CA due to its inherent randomness in providing the channel access cannot provide guarantees on airtime, throughput, or delay jitter, and it suffers significantly as the number of associated clients increases. Another issue with the probabilistic access arises with the hidden terminal problem [5], when clients are within range of the AP, but out of range with each other. It is tackled with the RTS/CTS mechanism which adds to the transmission overhead.
I-D Deterministic channel access
@@
We propose to use the Target Wake Time feature in Wi-Fi 6 to provide a predetermined channel access schedule, allowing transmission of packets at specific predefined intervals and show its effectiveness over legacy CSMA/CA based scheduling. To the best of our knowledge, our work is the first demonstration of this idea on an experimental WLAN testbed, composed of only consumer-grade off-the-shelf components.
I-E Contributions
@@
Organization: The paper is organized as follows. In Section II, we give an introduction to TWT. Section III provides an overview of related work. Sections IV and V describe, respectively, the experimental testbed and performance analysis of TWT over it to gain insights into working of TWT available on off-the-shelf components. Based on these insights, in Section VI, we formulate an optimization problem to find the optimal TWT-based deterministic schedule and system parameter values to maximize the overall throughput, and in Section VII, we offer a practical approach to solve this problem. Using the optimal values of various system parameters thus obtained for our WLAN testbed, in Section VIII we configure the testbed accordingly and experimentally compare the performance of the system with the proposed deterministic scheduling and the legacy random channel access. Our results show that deterministic scheduling not only improves the performance of individual TWT clients, but of the overall system as well. Finally, Section IX concludes the paper and provides some directions for the future work.
@@ III Related Work
Schneider et al. [11] demonstrate a proof of concept deterministic scheduling system for Industrial-IoT by using Wi-Fi 6 features of TWT and OFDMA, as an extension of the wired time-sensitive networks. A real time testbed is used to evaluate the jitter and latency, showing the potential of TWT for deterministic scheduling. In [12], the authors propose a combination of multi-link operation and TWT based optimization for energy conservation. The work discusses the trade-off of using Broadcast TWT, which saves energy but is inefficient in high traffic scenarios, and multi-link operation, which is favorable in high traffic situations but is energy hungry. The authors propose a joint optimization approach to save energy while meeting the traffic demands in a congested scenario. Hegde et al. [13] look into a practical implementation of downlink flow control where an intermediate proxy controls the TCP traffic by allocating time-slices to each user and setting up priorities by marking the WLAN packets differently. The work also controls the uplink TCP rate by controlling the flow of the ACKs. Sheth et al. [14] use individual TWT to provide non-overlapping schedules to schedule traffic for bursty, constant rate, and sparse periodic IoT sensor data, and compare metrics like channel occupancy and retransmission rate along with traditional metrics. The work in [15] looks into the latency requirements of commercial internet applications. The authors classify internet traffic into categories based on latency tolerance and use state machines to classify the applications based on packet arrival patterns in real time. The authors use an approach of matching the applications to the predefined TWT intervals and also update the cycles when deemed necessary. The work provides a practical testbed example of using deterministic scheduling to protect the applications from the effects of contention. Authors in [16] test the performance of TWT for synthetic video streaming traffic in a congested scenario. They only consider a specific traffic class with the traffic congestion created using only downlink traffic, and test the performance of a single TWT client. Puligheddu et al. [17] propose a TWT based scheduling algorithm for latency sensitive Industrial-IoT applications, using a scheduling strategy that respects the Age of Information (AoI) of the packets while reducing the net energy consumed, while demonstrating the effectiveness of TWT as a channel access mechanism in an industrial setting. However, the authors do not address the fundamental issue of congestion that largely plagues dense urban and enterprise Wi-Fi deployments and their interaction with the legacy network protocol stack and implementation issues in real life commodity hardware that severely affect the throughput of Wi-Fi. Authors in [18] highlight the need for using time sensitive networking and its superior performance when compared to legacy CSMA/CA based channel access, while also highlighting the need to simultaneously satisfy the need of high throughput and low latency often associated with demands of interactive AR/VR applications. The authors reckon that scheduled access as a combination of OFDMA, TWT combined with smarter MAC layer scheduling is the future of Wi-Fi based networks, even in traffic agnostics scenarios.
@@ VII The Proposed Solution approach
where the vectors $WT_{sample}\text{ and }ST_{sample}$ contain the corresponding WTs and STs, and $I_{i}$ is the indicator variable vector used to pick the values according to Eqn. (2).
$T_{i}=T_{sample,i}\times I_{i}^{\intercal},$ (2)
@@ VIII-B Buffer Bloat and Latency
When there are sudden large spikes in the packet traffic on the network, those cause the ingress buffers to be filled, resulting in increased latency experienced by the end users. This is usually experienced in the Wi-Fi scenarios as the Wi-Fi link is, generally, the bottle-neck link in the network. While QoS filters and EDCA do offer respite from this, it is still desirable to reduce such traffic “burstiness” by smoothening the incoming traffic. The graph in Figure 12 illustrates the smoothening of the bursts (3MB worth of data burst every 6 seconds) with TWT schedules. This smoothening could also allow the control of TCP congestion window, hence reducing the overall packet flow into the ingress buffers throughput the network and subsequently packet latency, by setting appropriate TWT cycles on the final Wi-Fi link, a effect similar to server side rate control mechanism suggested by the authors in [24].

### v2

@@ I Introduction
Carrier-Sense Multiple Access with Collision Avoidance (CSMA/CA) is a channel access method traditionally used by Wi-Fi devices. A CSMA/CA client listens to check for an idle channel before transmission. If the channel is indeed busy, it waits till the end of the transmission plus the Distributed Inter Frame Space (DIFS) duration before choosing a random delay. This delay is uniformly chosen from the contention window. If any other client transmits during this period, the countdown is paused and resumed after the transmission is completed. In case there is a collision, the client uses a binary exponential back-off mechanism before trying again, wherein the contention window is doubled, and a new delay is sampled from it. The back-off, on average, doubles after every failure, thus reducing the collision probability.
However, the CSMA/CA based channel access may cause significant penalties in terms of delay and throughput, while reducing the overall efficiency of the system due to airtime wasted because of the back-off mechanism. The CSMA/CA also impacts the fairness of the transmission. In a scenario where the traffic is downlink biased, yet a few clients with uplink traffic may receive a higher proportion of transmit time. This is because over a period of time the odds of every device winning the channel average out to be the same and when the Access Point (AP) wins the channel, its transmit time must be further divided among multiple downlink clients. This is often referred to as the uplink/downlink disparity of probabilistic channel access [4].
The CSMA/CA due to its inherent randomness in providing the channel access, cannot provide guarantees on airtime, throughput, or delay jitter, and it suffers significantly as the number of associated clients increases. Another issue with the probabilistic access arises with the hidden terminal problem [5], when clients are within range of the AP, but out of range with each other. It is tackled with the RTS/CTS mechanism which adds to the transmission overhead.
I-A Deterministic channel access
@@
We propose to use the Target Wake Time feature in Wi-Fi 6 to provide a predetermined channel access schedule, allowing transmission of packets at specific predefined intervals and show its effectiveness over legacy CSMA/CA based channel access. To the best of our knowledge, our work is the first demonstration of this idea on an experimental WLAN testbed, composed of only consumer-grade off-the-shelf components.
I-B Contributions
@@
I-C Organization
The paper is organized as follows. In Section II, we introduce TWT. Section III provides an overview of related work. Sections IV and V describe, respectively, the experimental testbed and performance analysis of TWT over it to gain insights into working of TWT available on off-the-shelf components. Based on these insights, in Section VI, we formulate an optimization problem to find the optimal TWT-based deterministic schedule and system parameter values to maximize the overall throughput, and in Section VII, we offer a practical approach to solve this problem. Using the optimal values of various system parameters thus obtained for our WLAN testbed, in Section VIII we configure the testbed accordingly and experimentally compare the performance of the system with the proposed deterministic scheduling and the legacy random channel access. Our results show that deterministic scheduling not only improves the performance of individual TWT clients, but of the overall system as well. Finally, Section IX concludes the paper and provides some directions for the future work.
@@ III Related Work
Schneider et al. [11] demonstrate a proof of concept deterministic scheduling system for Industrial-IoT by using Wi-Fi 6 features of TWT and OFDMA, as an extension of the wired time-sensitive networks. A real time testbed is used to evaluate the jitter and latency, showing the potential of TWT for deterministic scheduling. In [12], the authors propose a combination of multi-link operation and TWT based optimization for energy conservation. The work discusses the trade-off of using Broadcast TWT, which saves energy but is inefficient in high traffic scenarios, and multi-link operation, which is favorable in high traffic situations but is energy hungry. The authors propose a joint optimization approach to save energy while meeting the traffic demands in a congested scenario. Hegde et al. [13] look into a practical implementation of downlink flow control where an intermediate proxy controls the TCP traffic by allocating time-slices to each user and setting up priorities by marking the WLAN packets differently. The work also controls the uplink TCP rate by controlling the flow of the ACKs. Sheth et al. [14] use individual TWT to provide non-overlapping schedules to schedule traffic for bursty, constant rate, and sparse periodic IoT sensor data, and compare metrics like channel occupancy and retransmission rate along with traditional metrics. The work in [15] looks into the latency requirements of commercial internet applications. The authors classify internet traffic into categories based on latency tolerance and use state machines to classify the applications based on packet arrival patterns in real time. The authors use an approach of matching the applications to the predefined TWT intervals and also update the cycles when deemed necessary. The work provides a practical testbed example of using deterministic scheduling to protect the applications from the effects of contention. Authors in [16] test the performance of TWT for synthetic video streaming traffic in a congested scenario. They only consider a specific traffic class with the traffic congestion created using only downlink traffic, and test the performance of a single TWT client. Puligheddu et al. [17] propose a TWT based scheduling algorithm for latency sensitive Industrial-IoT applications, using a scheduling strategy that respects the Age of Information (AoI) of the packets while reducing the net energy consumed, while demonstrating the effectiveness of TWT as a channel access mechanism in an industrial setting. However, the authors do not address the fundamental issue of congestion that largely plagues dense urban and enterprise Wi-Fi deployments and their interaction with the legacy network protocol stack and implementation issues in real life commodity hardware that severely affect the throughput of Wi-Fi. Authors in [18] highlight the need for using time sensitive networking and its superior performance when compared to legacy CSMA/CA based channel access, while also highlighting the need to simultaneously satisfy the need of high throughput and low latency often associated with demands of interactive AR/VR applications. The authors reckon that scheduled access as a combination of OFDMA, TWT combined with smarter MAC layer scheduling is the future of Wi-Fi based networks, even in traffic agnostic scenarios.
@@ VII The Proposed Solution approach
where the vectors $WT_{sample}\text{ and }ST_{sample}$ contain the corresponding WTs and STs, and $I_{i}$ is the indicator variable vector used to pick the values according to the following set of equations:
$T_{i}=T_{sample,i}\times I_{i}^{\intercal},$
@@ VII-A Issues with this approach
Both issues can be addressed by one of the two approaches: including appropriate constraints in the above optimization problem or letting the optimization problem be as it is and applying the constraints corresponding to these two issues on its resulting solution. Technically, there is no difference between these two approaches, but the first approach, though it appears neat, is somewhat opaque in that it does not provide much insights into the application of the corresponding constraints and resulting solution, while the second approach lets us do so. Therefore, we adopt the second approach in the rest of the paper and introduce the concepts of “pseudo client” and “acceptable overlap” next to implement it.
@@ VIII-B Buffer Bloat and Latency
When there are sudden large spikes in the packet traffic on the network, causing the ingress buffers to be filled, resulting in increased latency experienced by the end users. This is usually experienced in the Wi-Fi scenarios as the Wi-Fi link is, generally, the bottle-neck link in the network. While QoS filters and EDCA do offer respite from this, it is still desirable to reduce such traffic “burstiness” by smoothening the incoming traffic. The graph in Figure 12 illustrates the smoothening of the bursts (3MB worth of data burst every 6 seconds) with TWT schedules. This smoothening could also allow the control of TCP congestion window, hence reducing the overall packet flow into the ingress buffers throughput the network and subsequently packet latency, by setting appropriate TWT cycles on the final Wi-Fi link, a effect similar to server side rate control mechanism suggested by the authors in [24].